a particle of mass m is projected with a velocity v making an angle of 45 degrees with the horizontal. the magnitude of the angular momentum of the projectile about the point pf projection when the particle is at its max height 

Grade:12

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Pranay

max height ,H= v2sin2θ/2g

velocity at max height = ux= vcosθ

angular momentum= muxH

and putting θ =450

angular momentum = mv3/g4√2

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

vikas askiitian expert
13 years ago

if the point of projection is taken as origen then

 

  coordinates of maximum height (H,R/2)       or r =Hi + (R/2)j       in vector  form

R is horizontal  range  and H is maximum height ....

 

at maximum height there is only horizontal component of velocity no vertical component exists ...

so , V= vcos@i  ......

   linear momentam = mvcos@i                     ( @ is the angle of projection)

 we know that angular momentam(L) about point of projection = r*p                (* is cross product)

                       L  =(Hi + (R/2)j)*(vi)

                       L  =Rv/2 (-K)

                     L=v3 /2g (-k)                                       ( R =v2 /g at @=45 degree)